OpenStudy (anonymous):

What number would you put in a matrix to represent a term that is missing in the original equation.? A. 1 B. 0 C. 2 D. -1

jimthompson5910 (jim_thompson5910):

if you had something like 2x+3z = 10 then that row will have 2, 0, 3, 10 the "0" represents 0y 2x+3z = 10 is the same as 2x+0y+3z = 10

OpenStudy (loser66):

What is "the original question?"